S000 for the data and calculate E (P) Exercise 9: Find a curve fit y P(t)1+ Ce 3 2 1 0 t 2800 1000 500 1800 y
Answer #1

ti, y){(0,500), (1,1000), (2, 1800), (3, 2800) is the data which we want to fit to the curve

5000 yP(t) 1+CeAt

Using the function fit functionality in Matlab, we get:

Command Window >> x=[0,1, 2,31; %xt y [500,1000,1800,2800]; FT-fittype (5000/ (1+C*exp (A*x)) ); %y=P (t) =P (x) f0-fit (x

And so we get the model as 5000 yP(t) 1+9.0282e-0.8135t

The data points together with the model curve approximation is given below:

X 3000 500 0 (3, 2800 1000 1 1800 2 2800 3 2000 (2.1800 5000 1+9,0282e-0.8135x 1000 (1,1000) (0,500 Label: 4 X 0 2 4 powered